OpenStudy (anonymous):

for what value of x is the rational expression below to zero x-8/x+8

OpenStudy (anonymous):

x=8

OpenStudy (anonymous):

how

OpenStudy (anonymous):

set it equal to 0

OpenStudy (anonymous):

\[\Large \frac{ x-8 }{ x+8 }=0\]

OpenStudy (anonymous):

multiply both sides by (x+8)

OpenStudy (anonymous):

\[\Large (x+8) \times \frac{ x-8 }{ x+8 }=0 \times (x+8)\]

OpenStudy (anonymous):

x-8=0

OpenStudy (anonymous):

add 8 to both sides

OpenStudy (anonymous):

x=8

OpenStudy (anonymous):

okay thank you @doulikepiecauseidont